# Advanced Technology, Chinese Academy of Sciences, Shenzhen, China

### v1

@@ Abstract
In medical imaging, 4D MRI enables dynamic 3D visualization, yet the trade-off between spatial and temporal resolution requires prolonged scan time that can compromise temporal fidelity—especially during rapid, large-amplitude motion. Traditional approaches typically rely on registration-based interpolation to generate intermediate frames. However, these methods struggle with large deformations, resulting in misregistration, artifacts, and diminished spatial consistency. To address these challenges, we propose TSSC-Net, a novel framework that generates intermediate frames while preserving spatial consistency. To improve temporal fidelity under fast motion, our diffusion-based temporal super-resolution network generates intermediate frames using the start and end frames as key references, achieving 6×6$\times$6 × temporal super-resolution in a single inference step. Additionally, we introduce a novel tri-directional Mamba-based module that leverages long-range contextual information to effectively resolve spatial inconsistencies arising from cross-slice misalignment, thereby enhancing volumetric coherence and correcting cross-slice errors. Extensive experiments were performed on the public ACDC cardiac MRI dataset and a real-world dynamic 4D knee joint dataset. The results demonstrate that TSSC-Net can generate high-resolution dynamic MRI from fast-motion data while preserving structural fidelity and spatial consistency.
@@ 1 Introduction
The evolution of four-dimensional magnetic resonance imaging (4D MRI) has revolutionized dynamic anatomical assessment, enabling volumetric visualization of physiological motions across time—a capability that is critical for applications such as cardiac function analysis [28, 21], tumor kinetics [13], and joint biomechanics investigations [8]. By acquiring time-resolved 3D image sequences, 4D MRI uniquely captures transient phenomena like myocardial contraction patterns and rapid synovial movements, providing clinicians with unprecedented insights into disease progression and treatment response. However, conventional 4D MRI acquisition protocols are hampered by inherent limitations [24, 29]. The fundamental trade-off between spatial and temporal resolution dictated by MRI physics necessitates prolonged scan durations, which in turn compromise the temporal fidelity of the acquired data [6, 25]. This limitation is more pronounced in cases involving rapid or large-amplitude anatomical motion. Consequently, the resulting sparse temporal sampling and potential spatial inconsistencies diminish the clinical utility of 4D MRI in time-sensitive diagnostic scenarios.
Recent advances in deep learning have spurred interest in data-driven approaches for 4D MRI generation. While generative adversarial networks (GANs) [9] and deformable registration techniques [1, 2, 5, 15] have demonstrated preliminary success in temporal interpolation, they exhibit three critical shortcomings: (1) GANs often exhibit inherent limitations in spatiotemporal sequence modeling, as their adversarial training mechanism fails to capture smooth temporal evolution, resulting in progressive distortion on temporal trajectories [3]. (2) Conventional registration-based interpolation methods often lack explicit mechanisms to enforce smooth inter-phase motion continuity, which may introduce structural distortions during the interpolation process [1, 2, 5, 15]. (3) Existing frameworks often lack explicit mechanisms to enforce cross-slice spatial consistency, which may lead to volumetric inconsistencies and reduce the reliability of subsequent 3D analyses [14, 11]. In contrast to GANs and registration-based methods, diffusion models [22, 20] demonstrate impressive generation performance in natural image synthesis but remain under-explored due to the unique coupled spatiotemporal challenge of MRI dynamics.
@@
    Our method delivers outstanding performance: achieving competitive results on ACDC cardiac MRI with minor deformations and significantly improving knee joint MRI with large motions. Notably, TSSC-Net can produce 6×6$\times$6 × temporal super-resolution in a single inference step.
@@ 2.2 Diffusion-driven temporal super-resolution
The forward diffusion process defines a Markov chain [12] that progressively adds noises to a clean image over $T$ steps. In TSSC-Net, we add noises to intermediate frames($x_{0}$) and train the model to predict the noises, conditioned on ${I_{0}}$ (start frame) and ${I_{1}}$ (end frame). This allows the model to learn the noise distribution of intermediate dynamics. Starting from $x_{0}$, the forward process generates a sequence ${x_{1},x_{2},…,x_{T}}$. For each step $t=1,…,T$,
@@
with ${\beta_{t}}{\epsilon}{(0,1)}$ being the noises variance at step $t$. This forward process is fixed (non-learnable) and is used to provide training pairs of noisy images and their corresponding clean target. In experiments, T is set to 1000.
@@
Furthermore, as shown in Stage 1 of Fig. 1, we employ a transformer-based [19] architecture within our diffusion model that partitions the noisy image into a sequence of tokens. By integrating a spatiotemporal attention mechanism, the model effectively captures cross-frame correlations in both spatial and temporal dimensions. This design not only enhances the denoising process but also ensures that the generated frames exhibit consistent structural details and realistic motion patterns. Finally, leveraging this cross-frame temporal attention enables robust 6×6$\times$6 × frame interpolation, yielding synthesized outputs with high visual quality and temporal coherence.
@@ 2.3 Spatial Consistency Enhancement
After the temporal super-resolution stage, the independently generated 2D slices are concatenated into initial 3D volumes, which may lead to cross-slice inconsistencies and discontinuities. To address these issues, we adopt the Mamba architecture, a framework built on state space models (SSMs) [10, 27] that overcomes limitations of conventional methods such as the quadratic complexity of transformers and the restricted receptive fields of CNNs. The Mamba architecture is particularly effective for processing 3D images and long sequences, as it efficiently captures long-range dependencies [17]. Specifically, the hidden state is updated and the output is generated via:
@@
where $h_{t}$ denotes the hidden state at time t, $x_{t}$ is the input, and $\bar{A}_{t}$, $\bar{B}_{t}$ and $\bar{C}_{t}$ are learned parameters. The framework is designed to capture long-range dependencies while maintaining computational efficiency. To address spatial inconsistencies in the initial 3D volumes, we introduce a spatial consistency enhancement network that employs residual tri-directional Mamba blocks. Each block processes the 3D volume along three orthogonal directions, denoted as $\overleftrightarrow{SSM_{xyz}}$, $\overleftrightarrow{SSM_{yzx}}$, and $\overleftrightarrow{SSM_{zxy}}$, to capture horizontal, vertical, and inter-slice dependencies, respectively. The bidirectional structure of each SSM, capturing both forward and backward dependencies along its respective axis, enables robust integration of multi-directional contextual information.
@@ 3.1 Dataset
To verify the proposed method for 4D image generation, two 4D image datasets are used, each for the heart and knee. The publicly available ACDC [4] dataset consists of 150 4D cardiac MRI scans, each spanning a full cardiac cycle from end-diastole (ED) to end-systole (ES). All scans were resampled in the x–y plane using bilinear interpolation, and zero-padding was applied along the z-axis to obtain a standardized volume size of 256×256×32. The dataset was split into 100 training samples and 50 testing samples. To assess generalizability, we also collected an in-house 4D dynamic knee MRI dataset (12 cases) using a United Imaging 3T scanner. Each scan consists of 12 3D frames covering a full flexion-extension cycle, acquired with a coronal T1-weighted sequence. The volumes were resampled to 256×256×96 to ensure consistent spatial resolution. In total, 1,152 2Dt slices were used for the first-stage temporal super-resolution task, and 144 3D volumes were used for the second-stage spatial consistency enhancement.
@@ 3.2 Implementation Details
Experiments were conducted using PyTorch 11.7 on an NVIDIA A100 80GB GPU. In the first stage, the diffusion-driven temporal super-resolution network was trained for 100,000 iterations using the Adam optimizer with a fixed learning rate of $1\times 10^{-4}$. A linear noise schedule was applied, with ${{\beta}_{t}}$ increasing from ${10^{-6}}$ to ${10^{-2}}$, and DDIM acceleration was employed during inference to expedite the sampling process. This stage focused on achieving temporal super-resolution by generating high-quality intermediate frames. Once the diffusion-driven temporal super-resolution network was fully trained, its parameters were frozen. In the second stage, the spatial consistency enhancement network was then trained for 100 epochs using the Adam optimizer. The training began with a learning rate of $1\times 10^{-4}$ that was linearly decayed over time, with the aim of refining the 3D voxel coherence and eliminating cross-slice inconsistencies.
@@ 3.3 Results and Discussion
In our experiments, TSSC-Net was benchmarked against three registration-based baselines: VM [2], TM [5], and UVI-Net [16]. These methods estimate flow fields from ${I_{0}}$ to ${I_{1}}$ and use linear interpolation to generate intermediate frames. VM (VoxelMorph) is a registration-based method that aligns images at the voxel level. TM (TransMorph) builds upon this by integrating transformer modules to capture global context and enhance registration accuracy. Conversely, UVI-Net is an unsupervised bidirectional registration interpolation approach that synthesizes intermediate frames through bidirectional flow estimation.
@@ 4 Conclusion
In this paper, we introduced TSSC-Net, a novel 4D MRI imaging framework that combines a diffusion-driven temporal super-resolution network with a spatial consistency enhancement network based on residual tri-directional Mamba blocks. Experiments on both the ACDC cardiac dataset and an in-house knee joint dataset demonstrate that TSSC-Net achieves competitive performance on low-motion data and significantly outperforms conventional registration-based methods in high-motion scenarios by mitigating cross-slice inconsistencies and preserving anatomical integrity. Notably, TSSC-Net can generate a complete 12-frame 4D dynamic MRI sequence from only two input frames in a single inference step, achieving 6×6$\times$6 × temporal super-resolution. These results underscore its potential for reliable 4D MRI imaging, and future work will focus on further enhancing robustness and adapting the framework to diverse anatomical regions and motion patterns. The code will be available at https://github.com/Joker-ZXR/TSSC-Net.
{credits}
4.0.1 Acknowledgements
This research was partly supported by the National Natural Science Foundation of China (No. 62222118, No. U22A2040), Shenzhen Science and Technology Program (No. JCYJ20220531100213029), Shenzhen Medical Research Fund (No. B2402047), National Key R&{\&}&D Program of China (No. 2023YFA1011400), Key Laboratory for Magnetic Resonance and Multimodality Imaging of Guangdong Province (No. 2023B1212060052), and Youth lnnovation Promotion Association CAS.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@ Abstract
In medical imaging, 4D MRI enables dynamic 3D visualization, yet the trade-off between spatial and temporal resolution requires prolonged scan time that can compromise temporal fidelity—especially during rapid, large-amplitude motion. Traditional approaches typically rely on registration-based interpolation to generate intermediate frames; however, these methods struggle with large deformations, resulting in misregistration, artifacts, and diminished spatial consistency. To address these challenges, we propose TSSC-Net, a novel framework that generates intermediate frames while preserving spatial consistency. To solve temporal fidelity under fast motion, our diffusion-based temporal super-resolution network generates intermediate frames using the start and end frames as key references, achieving 6× temporal super-resolution in a single inference step. Additionally, we introduce a novel tri-directional mamba-based module that leverages long-range contextual information to effectively resolve spatial inconsistencies arising from cross-slice misalignment, thereby enhancing volumetric coherence and correcting cross-slice errors. Extensive experiments were performed on the public ACDC cardiac MRI dataset and a real-world dynamic 4D knee joint dataset. The results demonstrate that TSSC-Net can generate high-resolution dynamic MRI from fast-motion data while preserving structural fidelity and spatial consistency. The code will be available at https://github.com/Joker-ZXR/TSSC-Net.
@@ 1 Introduction
The evolution of four-dimensional magnetic resonance imaging (4D MRI) has revolutionized dynamic anatomical assessment, enabling volumetric visualization of physiological motions across time—a capability that is critical for applications such as cardiac function analysis [25, 20], tumor kinetics [13], and joint biomechanics investigations [8]. By acquiring time-resolved 3D image sequences, 4D MRI uniquely captures transient phenomena like myocardial contraction patterns and rapid synovial movements, providing clinicians with unprecedented insights into disease progression and treatment response. However, conventional 4D MRI acquisition protocols are hampered by inherent limitations. The fundamental trade-off between spatial and temporal resolution dictated by MRI physics necessitates prolonged scan durations, which in turn compromise the temporal fidelity of the acquired data [6]. This is especially problematic in scenarios where organs or joints exhibit rapid or large-amplitude displacements. Consequently, the resulting sparse temporal sampling and potential spatial inconsistencies diminish the clinical utility of 4D MRI in time-sensitive diagnostic scenarios.
Recent advances in deep learning have spurred interest in data-driven approaches for 4D MRI generation. While generative adversarial networks (GANs) [9] and deformable registration techniques [1, 2, 5, 15] have demonstrated preliminary success in temporal interpolation, they exhibit three critical shortcomings: (1) GANs often exhibit inherent limitations in spatiotemporal sequence modeling, as their adversarial training mechanism fails to capture smooth temporal evolution, resulting in progressive distortion on temporal trajectories[3]; (2) Conventional registration-based interpolation methods often lack explicit mechanisms to enforce smooth inter-phase motion continuity, resulting in unnatural structural distortions during the interpolation process [1, 2, 5, 15]; (3) Existing frameworks neglect cross-slice spatial consistency, causing volumetric inconsistencies that invalidate 3D analyses[14, 11]. In contrast to GANs and registration-based methods, diffusion models[21, 19] demonstrate impressive generation performance in natural image synthesis but remain under-explored due to the unique coupled spatiotemporal challenge of MRI dynamics.
@@
    Our method delivers outstanding performance: achieving competitive results on ACDC cardiac MRI with minor deformations and significantly improving knee joint MRI with large motions. Notably, TSSC-Net can produce 6× temporal super-resolution in a single inference step.
@@ 2.2 Diffusion-driven temporal super-resolution
The forward diffusion process defines a Markov chain [12] that progressively adds noise to a clean image over $T$ steps. In TSSC-Net, this process is applied to the target frame ${I_{1}}$. Starting from $x_{0}=I_{1}$, the forward process generates a sequence ${x_{1},x_{2},…,x_{T}}$. For each step $t=1,…,T$ ,
@@
with ${\beta_{t}}{\epsilon}{(0,1)}$ being the noise variance at step $t$. This forward process is fixed (non-learnable) and is used to provide training pairs of noisy images and their corresponding clean target.
@@
Furthermore, as shown in Stage 1 of Fig. 1, we employ a transformer-based [18] architecture within our diffusion model that partitions the noisy image into a sequence of tokens. By integrating a spatiotemporal attention mechanism, the model effectively captures cross-frame correlations in both spatial and temporal dimensions. This de-sign not only enhances the denoising process but also ensures that the generated frames exhibit consistent structural details and realistic motion patterns. Finally, lev-eraging this cross-frame temporal attention enables robust 6× frame interpolation, yielding synthesized outputs with high visual quality and temporal coherence.
@@ 2.3 Spatial Consistency Enhancement
After the temporal super-resolution stage, the independently generated 2D slices are concatenated into initial 3D volumes, which can lead to cross-slice inconsistencies and discontinuities. To address these issues, we adopt the Mamba architecture, a framework built on state space models (SSMs) [10, 24] that overcomes limitations of conventional methods such as the quadratic complexity of transformers and the restricted receptive fields of CNNs. The Mamba architecture is particularly effective for processing 3D images and long sequences, as it efficiently captures long-range dependencies. Specifically, the hidden state is updated and the output is generated via:
@@
where $h_{t}$ denotes the hidden state at time $x_{t}$ is the input, and $\bar{A}_{t}$,$\bar{B}_{t}$ and $\bar{C}_{t}$ are learned parameters. This framework is highly advantageous as it effectively captures long-range dependencies with efficient computational scaling. To resolve the spatial inconsistencies, present in the initial 3D volumes, we propose a spatial consistency enhancement network that leverages residual tri-directional mamba blocks. Each block processes the 3D volume along three orthogonal directions, denoted as $\overleftrightarrow{SSM_{xyz}}$, $\overleftrightarrow{SSM_{yzx}}$, and $\overleftrightarrow{SSM_{zxy}}$, to capture horizontal, vertical, and inter-slice dependencies, respectively. The bidirectional structure of each SSM, capturing both forward and backward dependencies along its respective axis, enables robust integration of multi-directional contextual information.
@@ 3.1 Dataset
To verify the proposed method for 4D image generation, two 4D image datasets are used, each for the heart and knee. The publicly available ACDC [4] dataset consists of 150 4D cardiac MRI scans. To standardize the data, resampling was performed and zero-padding was applied along the z-axis, resulting in a final voxel volume of 256×256×32. The dataset was split into 100 training samples and 50 testing samples. To further evaluate generalizability, an in-house 4D dynamic knee MRI dataset (12 cases) was acquired using a United Imaging 3T scanner. The knee flexion angles ranged from 0° to 80°, and each 4D scan contained 12 frames of 3D MRI captured with a coronal T1-weighted sequence. The voxel size was resampled to 256×256×96 for consistency. In total, 1,152 2Dt slices were used for the first-stage temporal super-resolution task, and 144 3D volumes were used for the second-stage spatial consistency enhancement.
@@ 3.2 Implementation Details
Experiments were conducted using PyTorch 11.7 on an NVIDIA A100 GPU. In the first stage, the diffusion-driven temporal super-resolution network was trained for 100,000 iterations using the Adam optimizer with a fixed learning rate of $1\times 10^{-4}$. A linear noise schedule was applied, with ${{\beta}_{t}}$ increasing from ${10^{-6}}$ to ${10^{-2}}$, and DDIM acceleration was employed during inference to expedite the sampling process. This stage focused on achieving temporal super-resolution by generating high-quality intermediate frames. Once the diffusion-driven temporal super-resolution network was fully trained, its parameters were frozen. In the second stage, the spatial consistency enhancement network was then trained for 100 epochs using the Adam optimizer. The training began with a learning rate of $1\times 10^{-4}$ that was linearly decayed over time, with the aim of refining the 3D voxel coherence and eliminating cross-slice inconsistencies
@@ 3.3 Results and Discussion
In our experimental evaluation, the proposed TSSC-Net was benchmarked against three baseline methods: VM, TM, and UVI-Net. VM (VoxelMorph) [2] is a registration-based method that aligns images at the voxel level. TM (TransMorph) [5] builds upon this by integrating transformer modules to capture global context and enhance registration accuracy. Conversely, UVI-Net [16] is an unsupervised bidirectional registration interpolation approach that synthesizes intermediate frames through bidirectional flow estimation.
@@ 4 Conclusion
In this paper, we introduced TSSC-Net, a novel 4D MRI imaging framework that combines a diffusion-driven temporal super-resolution network with a spatial consistency enhancement network based on residual tri-directional mamba blocks. Experiments on both the ACDC cardiac dataset and an in-house knee joint dataset demonstrate that TSSC-Net achieves competitive performance on low-motion data and significantly outperforms conventional registration-based methods in high-motion scenarios by mitigating cross-slice inconsistencies and preserving anatomical integrity. Notably, TSSC-Net can generate a complete 12-frame 4D dynamic MRI sequence from only two input frames in a single inference step, achieving 6× temporal super-resolution. These results underscore its potential for reliable 4D MRI imaging, and future work will focus on further enhancing robustness and adapting the framework to diverse anatomical regions and motion patterns.